Determine the equation of the line through (-4,7) with a slope of -2. Write the answer in standard form using only integers as the coefficients.
Answer by tutorcecilia(2152)

Determine the equation of the line through (-4,7) with a slope of -2. Write the answer in standard form using only integers as the coefficients.
Use the point-slope formula:
(y-y1)=m(x-x1) [Plug-in the points]
(y-7) = (-2)(x-(-4))[Simplify]
y-7=-2(x+4)
y-7=-2x - 8 [Move the variables to one side of the equation]
y = -2x - 1
2x + y = - 1 [Standard form]